OpenStudy (onepieceftw):

PLEASE HELP ;_; Sandra is riding the Ferris wheel, and her height can be modeled by the equation H(t) = 25 cos (pi/14*t)+ 31, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Sandra before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Sandra begins the ride at the top, how far from the ground is the edge of the Ferris wheel when Sandra's height above the ground reaches a minimum?

OpenStudy (onepieceftw):

\[H(t)= 25 \cos (\frac{ \pi }{ 14 }t)+31\]